# Beyond Structure: Invariant Crystal Property Prediction with Pseudo-Particle Ray Diffraction

### v1

@@ Abstract
Crystal property prediction, governed by quantum mechanical principles, is computationally prohibitive to solve exactly for large many-body systems using traditional density functional theory. While machine learning models have emerged as efficient approximations for large-scale applications, their performance is strongly influenced by the choice of atomic representation. Although modern graph-based approaches have progressively incorporated more structural information, they often fail to capture long-term atomic interactions due to finite receptive fields and local encoding schemes. This limitation leads to distinct crystals being mapped to identical representations, hindering accurate property prediction. To address this, we introduce PRDNet that leverages unique reciprocal-space diffraction besides graph representations. To enhance sensitivity to elemental and environmental variations, we employ a data-driven pseudo-particle to generate a synthetic diffraction pattern. PRDNet ensures full invariance to crystallographic symmetries. Extensive experiments are conducted on Materials Project, JARVIS-DFT, and MatBench, demonstrating that the proposed model achieves state-of-the-art performance. The code will be released once our paper is accepted.
@@ 1 Introduction
Material properties are fundamentally determined by quantum mechanical equations such as the Schrödinger equation (Kohn et al., 1996), which are intractable for many-body systems and thus commonly approximated by Density Functional Theory (DFT). However, the high computational cost of DFT limits its application to large atomic systems. Recently, machine learning (ML) force fields have emerged as scalable surrogates (Chen et al., 2019; Chen & Ong, 2022), trained on DFT data to enable efficient structure relaxation and property prediction across diverse materials. These models achieve near-DFT accuracy while allowing exploration of large-scale atomic correlations.
In particular, the performance of ML models strongly depends on how atomic systems are represented (Xie & Grossman, 2018; Schütt et al., 2017; Chen et al., 2019; Louis et al., 2020; Choudhary & DeCost, 2021; Yan et al., 2022; 2024; Taniai et al., ; Ito et al., 2025). Early efforts relied on hand-engineered descriptors to capture atomic interactions (Isayev et al., 2017; Su et al., 2024; Zhu et al., 2024), which offered interpretability but suffered from human bias and poor transferability. With the emergence of data-driven models, Graph Neural Networks have become mainstream crystal encoding models. Specifically, they treat crystals as graphs and perform message passing (Chen et al., 2019) to aggregate node information. Initial solutions focus solely on pairwise local interactions, leveraging graph convolution networks (Xie & Grossman, 2018; Chen et al., 2019) and graph attention networks (Louis et al., 2020). To capture higher-order geometry, later works incorporated atomic angular embedding (e.g., bond angles and multi-body descriptors) (Choudhary & DeCost, 2021; Chen & Ong, 2022), improving structural fidelity but at the expense of efficiency and symmetry invariance. Most recently, equivariant/invariant GNNs (Yan et al., 2022; 2024; Taniai et al., ; Ito et al., 2025) have emerged as the leading paradigm, systematically overcoming prior limitations by embedding symmetry constraints such as periodicity-preserving message passing, thereby achieving state-of-the-art performance in Crystal Property Prediction (CPP).
@@
However, crystals are, in principle, infinite three-dimensional atomic systems. Current real-space encoders often fail to capture long-term atomic interactions (Tantivasadakarn et al., 2024), which are essential for determining physical properties. In particular, these methods typically struggle to represent them, as real-space embeddings with finite receptive fields in multi-edge graph constructions limit their ability to account for long-term effects. As a result, they map distinct crystal structures into the same representation. Figure 2 illustrates these examples. In DFT, these interactions are addressed through boundary conditions and supercell configurations (Makkar & Ghosh, 2021). A promising alternative is diffraction-space encoding: since every atom contributes to diffraction along specific lattice planes, ideal representation is guaranteed to embed complete real-space information without loss (Bouwkamp, 2016), enabling more faithful modeling of long-term correlations. Figure 1 illustrates the relationship between the crystal graph and its diffraction. Diffraction arises from the periodic arrangement of atoms and their interactions with the probing beam, described by form factors. Traditional methods, such as X-ray diffraction, rely on fixed, tabulated atomic form factors that depend only on the scattering vector and atomic species, deriving from historically collected, averaged quantum mechanical calculations and experimental measurements. This makes them incapable of distinguishing between atoms of the same element situated in different local chemical environments, as they ignore crucial structural variations that affect material properties.
@@
In this work, we leverage long-term atomic interactions by incorporating reciprocal-space (diffraction-based) information into crystal representations. Reciprocal space provides a compact, non-approximate description of crystal structures, naturally capturing long-term periodic interactions. To exploit this, we introduce a novel pseudo-particle representation, named PRDNet, that serves as a unique probe, learned through neural networks, and shown to be more effective for differentiating atoms and structural environments than conventional physical probes such as X-rays, electrons, or neutrons. Building on this idea, we develop a multimodal framework that integrates both the graph and diffraction information into the learning process. In addition, we show that the proposed model fulfills the geometric constraints, invariant to rotations, reflections, and translations of crystals. Quantitative evaluations on large-scale datasets, including Materials Project, JARVIS-DFT, and MatBench, demonstrate that our method consistently surpasses state-of-the-art models in common crystal property prediction tasks. In summary, our main contributions are as follows :
@@
  • We highlight the fundamental challenge of crystal representation, indicating that existing methods map distinct crystal structures to the same representation. To address this, we introduce diffraction-based representations, which is physically guaranteed to be unique.
@@ 2 Related work
Current crystal representations can be broadly categorized into three types: (1) multi-edge graph models; (2) atomic angular embedding; (3) periodic vector-based reference system. Despite their differences, their common limitation is mapping distinct crystal structures onto the same representations, as shown in Figure 2. Specifically, multi-edge crystal graphs, as in CGCNN (Xie & Grossman, 2018), SchNet (Schütt et al., 2017), MEGNet (Chen et al., 2019), and GATGNN (Louis et al., 2020), define atomic connectivity through cutoff radii, continuous filters, or message-passing and attention mechanisms. They are unable to distinguish distinct crystal systems, since angular information is not considered (Figure 2 a). To address this, models such as ALIGNN (Choudhary & DeCost, 2021) and M3GNet (Chen & Ong, 2022) incorporate bond-angle features, although restricted cutoffs may still yield identical angular relationships (Figure 2 b). Finally, multiple methods extend beyond lattice vectors by introducing periodic vectors, as in MatFormer (Yan et al., 2022) and iConformer/eConformer (Yan et al., 2024), or by employing Transformer-based frameworks, such as Crystalformer (Taniai et al., ) and Crystalframer (Ito et al., 2025), to capture broader correlations. However, these models remain limited by their local encodings, which often fail to differentiate structures or capture long-term interactions (Figure 2 c).
Recently, ReGNet (Nie et al., 2025) also introduced a reciprocal-space block to model long-range correlations. Nevertheless, it may produce inconsistent representations by violating the structure dependence of form factors, the details are discussed in Appendix A.2.
@@ 3 Preliminaries
Crystal structures can be described using a fundamental unit, such as a primitive cell, conventional unit cell, Bravais lattice, or asymmetric unit—together with its associated atoms. The unit is defined by a $3\times 3$ lattice matrix that specifies its geometry in real space. Formally, a crystal structure is represented as $\mathcal{M}=(\mathbf{A},\mathbf{P},\mathbf{L})$, where $\mathbf{A}$ denotes atomic types, $\mathbf{P}=[\mathbf{p}_{1},\dots,\mathbf{p}_{N}]\in\mathbb{R}^{3\times N}$ contains the fractional or Cartesian coordinates of the $N$ atoms, and $\mathbf{L}=[\bm{\ell}_{1},\bm{\ell}_{2},\bm{\ell}_{3}]\in\mathbb{R}^{3\times 3}$ defines the periodic lattice in real space. A crystal is an infinite periodic structure, expressed as
@@
Diffraction refers to the interaction of probe particles with periodic structures (Cao et al., ; Guo et al., 2025). In crystals, diffraction maps the real-space atomic arrangement into momentum (reciprocal) space, where each diffraction feature encodes structural information along specific periodic directions. Since diffraction is uniquely determined by the electron distribution for a given probe, and the electron distribution itself is governed by the potential field arising from the infinite atomic arrangement, diffraction provides a unique representation of the real-space crystal structure (Kohn et al., 1996). This process is governed by the structure factor (Greenfield et al., 1971), which describes how atoms scatter incident radiation (e.g., X-rays in XRD). Formally, diffraction is the Fourier transform of the electron density $\rho(\mathbf{r})$ in real space, giving
@@
where $\mathbf{r}$ is the Cartesian position vector and $\mathbf{Q}$ the basis on reciprocal-space. The electron density $\rho(\mathbf{r})$ reflects the spatial distribution of electrons and governs the dispersion properties of the crystal.
@@
Conventional diffraction probes, including X-ray photons, electrons, and neutrons, have intrinsic limitations in reciprocal-space representations (Bouwkamp, 2016). X-rays (Epp, 2016; Cao et al., ) and electrons (Glauber & Schomaker, 1953) scatter from the electron cloud, with intensities determined by the atomic form factor $f_{j}$, making atoms with similar charge nearly indistinguishable. Neutron diffraction (Shull & Wilkinson, 1953), in contrast, relies on nuclear scattering cross-sections, but heavy atoms of similar mass often show weak contrast. In a numerical framework, these limitations can be addressed by introducing a pseudo-particle with learned particle attributes, acting as a unique probe for reciprocal-space crystal representation, enabling differentiation of atomic arrangements in distinct local environments.
@@ 4.1 Overall Framework
This limitation motivates us to design a pseudo-particle that is explicitly sensitive to the local atomic environment while retaining strong discriminability across different elements, making it well-suited for CPP task. Unlike real particles, the learned pseudo-particle is not constrained by the physical limitations of X-rays. We define its form factor as:
@@
where $G_{\theta}(\mathcal{G}))$ encodes the local chemical environment learned by the graph network.
@@
A crystal structure is represented as a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where vertices $\mathcal{V}=\{v_{1},v_{2},\ldots,v_{N}\}$ correspond to atoms one-hot encoded attributes  (Xie & Grossman, 2018) and edges $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ encode atomic interactions within a cutoff radius. Each atom $i$ is characterized by a node feature $\mathbf{h}_{i}^{(0)}=\text{Embed}(v_{i})\in\mathbb{R}^{d}$ and coordinates $\mathbf{r}_{i}=(x_{i},y_{i},z_{i})\in\mathbb{R}^{3}$. Edge features between atoms $i$ and $j$ are defined as:
@@
Figure 3: The configuration of PRDNet, which integrates crystal attention and pseudo-particle diffraction to capture both short- and long-term atomic interactions.
@@
Crystals are inherently symmetrical. Following existing work (Yan et al., 2024), we force the model prediction to be invariant to $E(3)$ group (i.e., the rotation, reflection, and translation of input crystals). Specifically, according to Eq. 7, the final representation output $\mathbf{z}_{\text{fused}}$ depends on the $\mathbf{g}$ and $\mathbf{d}$. Since $\mathbf{g}$ depends on the relative geometric features $d_{ij}$ and $\theta_{ijk}$ (Eq.6) that remain unchanged under $E(3)$ operations, $\mathbf{g}$ is $E(3)$-invariant. Consequently, because $\mathbf{d}$ is also $E(3)$-invariant, the model output is $E(3)$-invariant, as shown in Section 4.3.

### v2

@@ Abstract
Crystal property prediction, governed by quantum mechanical principles, is computationally prohibitive to solve exactly for large many-body systems using traditional density functional theory. While machine learning models have emerged as efficient approximations for large-scale applications, their performance is strongly influenced by the choice of atomic representation. Although modern graph-based approaches have progressively incorporated more structural information, they often fail to capture long-range atomic interactions due to finite receptive fields and local encoding schemes. This limitation leads to distinct crystals being mapped to identical representations, hindering accurate property prediction. To address this, we introduce PRDNet that leverages unique reciprocal-space diffraction besides graph representations. To enhance sensitivity to elemental and environmental variations, we employ a data-driven pseudo-particle to generate a synthetic diffraction pattern. PRDNet ensures full invariance to crystallographic symmetries. Extensive experiments are conducted on Materials Project, JARVIS-DFT, and MatBench, demonstrating that the proposed model achieves state-of-the-art performance. The code is openly available at https://github.com/Bin-Cao/PRDNet.
@@ 1 Introduction
Material properties are fundamentally determined by quantum mechanical equations such as the Schrödinger equation (Kohn et al., 1996), which are intractable for many-body systems and thus commonly approximated by Density Functional Theory (DFT). However, the high computational cost of DFT limits its application to large atomic systems. Recently, machine learning (ML) force fields have emerged as scalable surrogates (Chen et al., 2019; Chen and Ong, 2022), trained on DFT data to enable efficient structure relaxation and property prediction across diverse materials. These models achieve near-DFT accuracy while allowing exploration of large-scale atomic correlations.
@@
In particular, the performance of ML models strongly depends on how atomic systems are represented (Xie and Grossman, 2018; Schütt et al., 2017; Chen et al., 2019; Louis et al., 2020; Choudhary and DeCost, 2021; Yan et al., 2022; 2024; Taniai et al., ; Ito et al., 2025). Early efforts relied on hand-engineered descriptors to capture atomic interactions (Isayev et al., 2017; Su et al., 2024; Zhu et al., 2024), which offered interpretability but suffered from human bias and poor transferability. With the emergence of data-driven models, Graph Neural Networks have become mainstream crystal encoding models. Specifically, they treat crystals as graphs and perform message passing (Chen et al., 2019) to aggregate node information. Initial solutions focus solely on pairwise local interactions, leveraging graph convolution networks (Xie and Grossman, 2018; Chen et al., 2019) and graph attention networks (Louis et al., 2020). To capture higher-order geometry, later works incorporated atomic angular embedding (e.g., bond angles and multi-body descriptors) (Choudhary and DeCost, 2021; Chen and Ong, 2022), improving structural fidelity but at the expense of efficiency and symmetry invariance. Most recently, equivariant/invariant GNNs (Yan et al., 2022; 2024; Taniai et al., ; Ito et al., 2025) have emerged as the leading paradigm, systematically overcoming prior limitations by embedding symmetry constraints such as periodicity-preserving message passing, thereby achieving state-of-the-art performance in Crystal Property Prediction (CPP).
@@
However, crystals are, in principle, infinite three-dimensional atomic systems. Current real-space encoders often fail to capture long-range atomic interactions (Tantivasadakarn et al., 2024), which are essential for determining physical properties. In particular, these methods typically struggle to represent them, as real-space embeddings with finite receptive fields in multi-edge graph constructions limit their ability to account for long-range effects. As a result, they map distinct crystal structures into the same representation. Figure 2 illustrates these examples. In DFT, these interactions are addressed through boundary conditions and supercell configurations (Makkar and Ghosh, 2021). A promising alternative is diffraction-space encoding: since every atom contributes to diffraction along specific lattice planes, ideal representation is guaranteed to embed complete real-space information without loss (Bouwkamp, 2016), enabling more faithful modeling of long-range correlations. Figure 1 illustrates the relationship between the crystal graph and its diffraction. Diffraction arises from the periodic arrangement of atoms and their interactions with the probing beam, described by form factors. Traditional methods, such as X-ray diffraction, rely on fixed, tabulated atomic form factors that depend only on the scattering vector and atomic species, deriving from historically collected, averaged quantum mechanical calculations and experimental measurements. This makes them incapable of distinguishing between atoms of the same element situated in different local chemical environments, as they ignore crucial structural variations that affect material properties.
Reciprocal-space representation is an effective strategy for capturing long-range interactions while remaining compact. This is because the complete diffraction pattern can be analytically derived from a single real-space unit cell, without requiring large supercell constructions, due to the intrinsic periodicity of crystals. However, achieving a lossless and physically valid reciprocal-space representation requires addressing several key gaps that existing works (Kosmala et al., 2023; Lin et al., 2023; Nie et al., 2025) have not systematically resolved. (1) The diffraction result must strictly satisfy the invariance constraints required by crystallographic symmetry. (2) Beyond real physical particles, which are inherently limited by physical laws and therefore not ideal as universal probes, a more flexible and learnable probing mechanism is needed. (3) The form factor must preserve its full physical dependencies, namely atomic species, local chemical environment, and diffraction-vector dependence, which are often simplified or omitted in prior methods. In addition, diffraction is fundamentally a global property of the entire structure, and therefore must be integrated at the modality level rather than treated as atom-wise feature fusion.
@@
To exploit this, we introduce a novel pseudo-particle representation, named PRDNet, that serves as a unique probe, learned through neural networks, and shown to be more effective for differentiating atoms and structural environments than conventional physical probes such as X-rays, electrons, or neutrons. Building on this idea, we develop a multimodal framework that integrates both the graph and diffraction information into the learning process. In addition, we show that the proposed model fulfills the geometric constraints, invariant to rotations, reflections, and translations of crystals. Quantitative evaluations on large-scale datasets, including Materials Project, JARVIS-DFT, and MatBench, demonstrate that our method consistently surpasses state-of-the-art models in common crystal property prediction tasks. In summary, our main contributions are as follows :
@@
  • We highlight the fundamental challenge of crystal representation, noting that existing methods can map distinct crystal structures to the same representation, and clarify the previously omitted premise of complete reciprocal-space information. To address this, we introduce diffraction-based representations, which are physically guaranteed to be unique.
@@
Figure 3: Analogy to light diffraction demonstrating how reciprocal space encodes long‐range interactions.
@@ 2 Related work
Current crystal representations can be broadly categorized into three types: (1) multi-edge graph models; (2) atomic angular embedding; (3) periodic vector-based reference system. Despite their differences, they share a common limitation: distinct crystal structures may be mapped onto identical representations, as shown in Figure 2. Specifically, multi-edge crystal graphs, as in CGCNN (Xie and Grossman, 2018), SchNet (Schütt et al., 2017), MEGNet (Chen et al., 2019), and GATGNN (Louis et al., 2020), define atomic connectivity through cutoff radii, continuous filters, or message-passing and attention mechanisms. They are unable to distinguish distinct crystal systems, since angular information is not considered (Figure 2 a). To address this, models such as ALIGNN (Choudhary and DeCost, 2021) and M3GNet (Chen and Ong, 2022) incorporate bond-angle features, although restricted cutoffs may still yield identical angular relationships (Figure 2 b). Finally, multiple methods extend beyond lattice vectors by introducing periodic vectors, as in MatFormer (Yan et al., 2022) and iConformer/eConformer (Yan et al., 2024), or by employing Transformer-based frameworks, such as Crystalformer (Taniai et al., ) and Crystalframer (Ito et al., 2025), to capture broader correlations (Figure 2 c). These approaches can satisfy basic crystallographic invariances, but they may still map distinct crystals to the same local neighborhood, particularly when the differences are governed by long-range periodicity or subtle symmetry variations.
Recent studies have highlighted the importance of long-range correlations in crystal representations and have sought to capture them through diffraction- or reciprocal-space descriptions. The early attempts aimed to introduce long-range descriptors in reciprocal space to capture the long-range interactions in Coulombic and van der Waals potentials, that are often omitted in ML force fields (Yu et al., 2022). The framework is straightforward: they apply Ewald summation to construct the reciprocal-space potential, which is then concatenated with real-space features to produce more comprehensive predictions. Following this line of work, models such as EwaldMP (Kosmala et al., 2023), PotNet (Lin et al., 2023), and ReGNet (renamed ReciNet in the latest version) (Nie et al., 2025) integrate Ewald summation/Fourier transform into message passing to model long-range interactions. However, existing methods often treat Ewald summation as a Fourier-like information fusion and overlook the unique and invariant nature of form factors, which should not be propagated through information aggregation across blocks or layers. Physically, form factors are functional quantities determined solely by the crystal structure and probe; they are unique, determined, and invariant.
In this work, we construct a complete invariant diffraction representation based on the fusion of structure and spectrum modalities, rather than atomic interaction information infusion, to explicitly encode long-range interactions. The diffraction-space description reflects the physical process of probing periodic atomic structures with incident particles. As illustrated by the analogy in Figure 3, the diffraction intensity at a reciprocal vector $\mathbf{Q}$ reflects the aggregated contribution of the entire real-space structure along different diffraction frequencies. That is, the response of an infinitely periodic atomic arrangement along that direction in crystal diffraction. By aggregating neighboring reciprocal vectors in a close group, we obtain multiple directional slices of the structure, which together provide a compact and physically consistent description of long-range interactions. A complete reciprocal-space representation depends on three key factors: (1) periodic atomic types $f_{i}^{\text{type}}$, (2) diffraction basis $|\mathbf{Q}|$, and (3) atomic local environments $G_{\theta}(\mathcal{G})$ (representing charge-density distributions), as detailed in Section 4.3. Existing models such as EwaldMP and PotNet approximate infinite-range potentials but ignore $G_{\theta}(\mathcal{G})$ in EwaldMP and both $G_{\theta}(\mathcal{G})$ and $|\mathbf{Q}|$ dependencies in PotNet, while ReGNet similarly applies a Fourier-like transformation in each reciprocal block without considering dependencies on $G_{\theta}(\mathcal{G})$ and $|\mathbf{Q}|$. Such omissions can lead to inconsistent reciprocal representations by violating the structure dependence and invariant of form factors, as discussed in Appendix A.3.
@@ 3 Preliminaries
Crystal structures can be described using a fundamental unit, such as a primitive cell, conventional unit cell, Bravais lattice, or asymmetric unit, together with its associated atoms. The unit is defined by a $3\times 3$ lattice matrix that specifies its geometry in real space. Formally, a crystal structure is represented as $\mathcal{M}=(\mathbf{A},\mathbf{P},\mathbf{L})$, where $\mathbf{A}$ denotes atomic types, $\mathbf{P}=[\mathbf{p}_{1},\dots,\mathbf{p}_{N}]\in\mathbb{R}^{3\times N}$ contains the fractional or Cartesian coordinates of the $N$ atoms, and $\mathbf{L}=[\bm{\ell}_{1},\bm{\ell}_{2},\bm{\ell}_{3}]\in\mathbb{R}^{3\times 3}$ defines the periodic lattice in real space. A crystal is an infinite periodic structure, expressed as
@@
Diffraction refers to the interaction of probe particles with periodic structures (Cao et al., ; Guo et al., 2025). In crystals, diffraction maps the real-space atomic arrangement into momentum (reciprocal) space, where each diffraction feature encodes structural information along specific periodic directions. Since diffraction is uniquely determined by the electron distribution for a given probe, and the electron distribution itself is governed by the potential field arising from the infinite atomic arrangement, diffraction provides a unique representation of the real-space crystal structure (Kohn et al., 1996; Cao et al., 2025). This process is governed by the structure factor (Greenfield et al., 1971), which describes how atoms scatter incident radiation (e.g., X-rays in XRD). Formally, diffraction is the Fourier transform of the electron density $\rho(\mathbf{r})$ in real space, giving
@@
where $\mathbf{r}$ is the Cartesian position vector and $\mathbf{Q}$ the basis in the reciprocal-space. The electron density $\rho(\mathbf{r})$ reflects the spatial distribution of electrons and governs the dispersion properties of the crystal.
@@
Unlike the Fourier transform recently applied for value extension, latent-space signal processing, or denoising, which decomposes an arbitrary signal into a frequency components to enrich the representation (Jiao et al., 2023; Choromanski et al., 2024; Dong et al., 2024; Gao et al., 2024), the Fourier transform in crystal diffraction is a physical scattering process. It maps the periodic electron density onto a discrete reciprocal lattice and provides a reciprocal view of the crystal structure under a different basis $\mathbf{Q}$, which physically reflects the crystal’s directional arrangement rather than its real-space Cartesian coordinates. Although the mathematical foundation is similar, the diffraction transform represents the accumulated phase differences arising from the periodic distribution/arrangement of electrons/atoms. The former yields a general frequency distribution, whereas the latter produces Bragg intensities that are strictly constrained by lattice periodicity and serve as a physical representation of the underlying crystal structure.
@@
Conventional diffraction probes, including X-ray photons, electrons, and neutrons, have intrinsic limitations in reciprocal-space representations (Bouwkamp, 2016). X-rays (Epp, 2016; Cao et al., ) and electrons (Glauber and Schomaker, 1953) scatter from the electron cloud, with intensities determined by the atomic form factor $f_{j}$, making atoms with similar charge nearly indistinguishable. Neutron diffraction (Shull and Wilkinson, 1953), in contrast, relies on nuclear scattering cross-sections, but heavy atoms of similar mass often show weak contrast. In a numerical framework, these limitations can be addressed by introducing a pseudo-particle with learned particle attributes, acting as a unique probe for reciprocal-space crystal representation, enabling differentiation of atomic arrangements in distinct local environments.
@@ 4.1 Overall Framework
This limitation motivates us to design a pseudo-particle that is explicitly sensitive to the local atomic environment while retaining strong discriminability across different elements, making it well-suited for CPP tasks. Unlike real particles, the learned pseudo-particle is not constrained by the physical limitations of X-rays. We define its form factor as:
@@
where $G_{\theta}(\mathcal{G})$ encodes the local chemical environment learned by the graph network.
@@
A crystal structure is represented as a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where vertices $\mathcal{V}=\{v_{1},v_{2},\ldots,v_{N}\}$ correspond to atoms one-hot encoded attributes  (Xie and Grossman, 2018) and edges $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ encode atomic interactions within a cutoff radius. Each atom $i$ is characterized by a node feature $\mathbf{h}_{i}^{(0)}=\text{Embed}(v_{i})\in\mathbb{R}^{d}$ and coordinates $\mathbf{r}_{i}=(x_{i},y_{i},z_{i})\in\mathbb{R}^{3}$. Edge features between atoms $i$ and $j$ are defined as:
@@
Figure 4: The configuration of PRDNet, which integrates crystal attention and pseudo-particle diffraction to capture both short- and long-range atomic interactions.
@@
Crystals are inherently symmetrical. Following existing work (Yan et al., 2024; Liu et al., 2024; Zheng et al., 2024), we force the model prediction to be invariant to $E(3)$ group (i.e., the rotation, reflection, and translation of input crystals). Specifically, according to Eq. 7, the final representation output $\mathbf{z}_{\text{fused}}$ depends on the $\mathbf{g}$ and $\mathbf{d}$. Since $\mathbf{g}$ depends on the relative geometric features $d_{ij}$ and $\theta_{ijk}$ (Eq.6) that remain unchanged under $E(3)$ operations, $\mathbf{g}$ is $E(3)$-invariant. Consequently, because $\mathbf{d}$ is also $E(3)$-invariant, the model output is $E(3)$-invariant, as shown in Section 4.3.